# A Linear Representation for Functions on Finite Sets

### v1

@@ Abstract
We demonstrate that any function $f$ from a finite set $Y$ to itself can be represented linearly. Specifically, we prove the existence of an injective map $j$ from $Y$ into a modular ring $\mathbb{Z}/m\mathbb{Z}$ and a constant $a\in\mathbb{Z}/m\mathbb{Z}$ such that the relation $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ holds for all $y\in Y$. This result is established by analyzing the algebraic properties of the adjugate of the characteristic matrix associated with the function’s digraph. The proof is constructive, providing a method for finding the embedding $j$, the modulus $m$, and the linear multiplier $a$.
Preprint Notice. This is an extended preprint version of a manuscript submitted to Journal of Combinatorial Theory, Series A. Licensed under CC BY-NC-ND 4.0. © 2025 Roman Bacik.
Keywords: Functional graph, Linear representation, Adjugate matrix, Characteristic polynomial, Discrete dynamical systems, Finite functions.
@@ 1 Introduction
This article establishes that any such function becomes linear after a suitable embedding. We show that for any function $f$ on a finite set, one can find an injective map $j$ from the set into a ring of integers modulo $m$ such that the action of $f$ corresponds to multiplication by a constant $a$. That is, the following relation holds for all $y\in Y$: $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ that the following diagram commutes:
@@
This result bridges the gap between arbitrary discrete maps and structured linear dynamical systems.
The formal proof of this result is under development in [5] using the Lean4. Without loss of generality, we can identify elements of the finite set $Y$ of size $n$ with the set $\{0,1,\ldots,n-1\}$, which are elements of $\mathbb{Z}/n\mathbb{Z}$.
@@ 2 Main Results
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. The function $f$ is represented by an $n\times n$ adjacency matrix $A=A_{f}$, where the entry $A_{ij}=\delta_{f(i),j}$ where $\delta_{i,j}$ is the Kronecker delta. With this convention, each row of $A$ contains exactly one non-zero entry.
@@
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Then $(A\cdot y)_{i}=y_{f(i)}$ for all $i\in\mathbb{Z}/n\mathbb{Z}$ and $y\in\mathbb{Z}^{n}$.
@@
2.1The proof follows from the Definition 2.1.
$(A\cdot y)_{i}=\sum_{j=0}^{n-1}A_{ij}y_{j}=\sum_{j=0}^{n-1}\delta_{f(i),j}y_{j}=y_{f(i)}$
@@
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Let $v\in\mathbb{Z}^{n}$ and $y=\operatorname{adj}(x\cdot I-A)\cdot v$. Then $y_{f(i)}=x\cdot y_{i}-m\cdot v_{i}$ where $m=\det(x\cdot I-A)$.
Proof.
2.2For adjugate matrix we have identity $(x\cdot I-A)\cdot\operatorname{adj}(x\cdot I-A)=\det(x\cdot I-A)\cdot I$. Therefore,
$m\cdot v=(x\cdot I-A)\cdot\operatorname{adj}(x\cdot I-A)\cdot v=(x\cdot I-A)\cdot y=x\cdot y-A\cdot y$
. The final equality follows from the Lemma 2.2. ∎
Lemma 2.4 (Determinant Degree Bound).
Let $M$ be an $n\times n$ matrix with polynomial entries $M_{ij}\in\mathbb{Z}[x]$. Then $\deg(\det(M))\leq\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}\deg(M_{ij})$.
@@
Lemma 2.5 (Polynomial Degree Properties of Adjugate Entries).
2.12.4Let $M=M(x)=\operatorname{adj}(x\cdot I-A)$ be the adjugate of the characteristic matrix $x\cdot I-A$. Then $M_{ij}=p_{ij}(x)$ is a polynomial in $x$ for all $i,j\in\mathbb{Z}/n\mathbb{Z}$ such that $p_{ii}(x)$ is monic of degree n1n-1 for all $i\in\mathbb{Z}/n\mathbb{Z}$ and $p_{ij}(x)$ has degree at most n2n-2 for all $i\neq j\in\mathbb{Z}/n\mathbb{Z}$.
@@
2.4The diagonal entries of $\operatorname{adj}(x\cdot I-A)$ are characteristic polynomials of $(n-1)\times(n-1)$ submatrices, hence monic of degree n1n-1. The off-diagonal entries correspond to minors with two diagonal positions removed, giving degree at most n2n-2 by Lemma 2.4. ∎
Lemma 2.6 (Polynomial with Positive Leading Coefficient - Integer Version).
If $p\in\mathbb{Z}[x]$ has positive leading coefficient, then $p(n)>0$ for all sufficiently large integers $n$.
@@
Write $p(n)=an^{d}+r(n)$ with $a=\text{leadingCoeff}(p)>0$, $d=\deg(p)$, and $\deg(r)<d$. For $n\geq 1$, we have $|r(n)|\leq B\cdot n^{d-1}$ where $B=\sum_{k<d}|p_{k}|$ (using integer arithmetic). Choose $n_{0}=\max(1,B)+1$. For $n>n_{0}$:
$p(n)=an^{d}+r(n)\geq an^{d}-Bn^{d-1}=n^{d-1}(an-B)>0$
since $a\geq 1$ (as $a$ is a positive integer) and $n>B$, so $an\geq n>B$. ∎
Lemma 2.7 (Strict Ordering of Adjugate-Vector Product).
2.52.6Let $M=M(x)=\operatorname{adj}(x\cdot I-A)$ be the adjugate of the characteristic matrix $x\cdot I-A$. Let $v=(0,1,2,\ldots,n-1)^{T}$. Then there is $x_{0}$ such that for all $x>x_{0}$:
  • $m(x)=\det(x\cdot I-A)>0$
  • The entries $y_{i}=(M(x)v)_{i}$ satisfy $y_{i}\geq 0$ for all $i$
  • The entries are strictly increasing: $y_{i}<y_{j}$ for all $i<j$
  • The entries are bounded by the determinant: $y_{i}<m(x)$ for all $i$
@@
2.52.6The proof follows from Lemma 2.5 and Lemma 2.6. Let $y=M(x)v$. Then $y_{i}=\sum_{k=0}^{n-1}M_{ik}(x)\cdot k$.
For each entry $y_{i}$, we express it as evaluation of a polynomial $p_{i}(x)=\sum_{k=0}^{n-1}M_{ik}(x)\cdot k\in\mathbb{Z}[x]$. By Lemma 2.5, the diagonal entry $M_{ii}$ is monic of degree n1n-1, while off-diagonal entries $M_{ik}$ (for $k\neq i$) have degree at most n2n-2. Therefore, the leading coefficient of $p_{i}$ is non-negative (dominated by the $M_{ii}\cdot i$ term with $i\geq 0$).
For the difference $p_{j}-p_{i}$ with $j>i$, the leading term comes from $(M_{jj}\cdot j-M_{ii}\cdot i)$. Since both $M_{jj}$ and $M_{ii}$ are monic of degree n1n-1, the leading coefficient of $p_{j}-p_{i}$ is ji>0j-i>0.
Similarly, for $p_{m}(x)=\det(x\cdot I-A)-p_{i}(x)$, since $\det(x\cdot I-A)$ is monic of degree $n$ (characteristic polynomial) and $p_{i}$ has degree at most n1n-1, the leading coefficient is 1.
@@
Definition 2.8 (Conversion from $\mathbb{Z}/n\mathbb{Z}$ to Index Set).
For computational purposes, we identify $\mathbb{Z}/n\mathbb{Z}$ with the index set $\{0,1,\ldots,n-1\}$ via the canonical map taking each residue class to its unique representative in $\{0,1,\ldots,n-1\}$.
@@
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. A linear representation of $f$ is an injective function $j:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/m\mathbb{Z}$ such that for all $i\in\mathbb{Z}/n\mathbb{Z}$,
@@
Theorem 2.10 (Main Theorem: Every Function Has Linear Representation).
2.12.82.32.72.9Any function $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ has a linear representation.
@@
2.82.32.7Let us define $m=\det(x\cdot I-A)$, $a=x$ and $j(i)=y_{i}$, where $y=M(x)v$ with $v=(0,1,2,\ldots,n-1)^{T}$. By Lemma 2.7, we can choose $x$ large enough such that:
  • $m>0$
  • The entries $y_{i}\geq 0$ are non-negative
  • The entries are strictly increasing: $y_{i}<y_{j}$ for $i<j$
  • The entries are bounded: $y_{i}<m$ for all $i$
The non-negativity and boundedness ensure that each $y_{i}\in[0,m)$ as an integer. The strict ordering property ensures that the map $j:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/m\mathbb{Z}$ defined by $j(i)=y_{i}\bmod m$ is injective.
Lemma 2.3 shows that $y_{f(i)}=x\cdot y_{i}-m\cdot v_{i}$. Since $v_{i}=i$ (using Definition 2.8), we have:
$j(f(i))=y_{f(i)}\bmod m=(x\cdot y_{i}-m\cdot i)\bmod m=x\cdot y_{i}\bmod m=x\cdot j(i)\bmod m$
Therefore, $j$ is a linear representation of $f$ with modulus $m$ and constant $a=x$. ∎
@@
Remark 2.12 (Explicit Computation).
The adjacency matrix for the quadratic function in Example 2.11 is:
@@
$x\cdot I-A_{f}=\begin{pmatrix}x-1&0&0\\ 0&x-1&0\\ 0&-1&x\end{pmatrix}$
@@
$\operatorname{adj}(x\cdot I-A_{f})=\begin{pmatrix}x(x-1)&0&0\\ 0&x(x-1)&0\\ 0&(x-1)&(x-1)^{2}\end{pmatrix}$
Using vector $v=(0,1,2)^{T}$, we get:
$y=\operatorname{adj}(x\cdot I-A_{f})\cdot v=\begin{pmatrix}x(x-1)\cdot 0\\ x(x-1)\cdot 1\\ (x-1)\cdot 1+(x-1)^{2}\cdot 2\end{pmatrix}=\begin{pmatrix}0\\ x^{2}-x\\ (x-1)(2x-1)\end{pmatrix}=\begin{pmatrix}0\\ x^{2}-x\\ 2x^{2}-3x+1\end{pmatrix}$
@@
$j(0)=0,\quad j(1)=12,\quad j(2)=21$
These values are strictly increasing ($0<12<21$) and bounded by $m=36$, so $j$ is injective.
@@
The lemma states that $y_{f(i)}=x\cdot y_{i}-m\cdot v_{i}$, which we can rewrite as:
$j(f(i))\equiv x\cdot j(i)\pmod{m}$
since the $m\cdot v_{i}$ term vanishes modulo $m$.
@@
$j(f(0))=j(0)=0$ $\equiv 4\cdot 0=0\pmod{36}\quad\checkmark$
$j(f(1))=j(1)=12$ $\equiv 4\cdot 12=48\equiv 12\pmod{36}\quad\checkmark$
$j(f(2))=j(1)=12$ $\equiv 4\cdot 21=84\equiv 12\pmod{36}\quad\checkmark$
Indeed: $48=36+12$, so $48\equiv 12\pmod{36}$, and $84=2\cdot 36+12$, so $84\equiv 12\pmod{36}$.

### v4

@@ Abstract
We demonstrate that any function $f$ from a finite set $Y$ to itself can be represented linearly. Specifically, we prove the existence of an injective map $j$ from $Y$ into a modular ring $\mathbb{Z}/m\mathbb{Z}$ and a constant $a\in\mathbb{Z}/m\mathbb{Z}$ such that $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ holds for all $y\in Y$. This result is established by analyzing the algebraic properties of the adjugate of the characteristic matrix associated with the function’s digraph. The proof is constructive, providing a method for finding the embedding $j$, the modulus $m$, and the linear multiplier $a$.
Keywords: Functional graph, Linear representation, Adjugate matrix, Characteristic polynomial, Finite functions.
@@ 1 Introduction
This article establishes that any such function becomes linear after a suitable embedding. We show that for any function $f$ on a finite set, one can find an injective map $j$ from the set into a ring of integers modulo $m$ such that the action of $f$ corresponds to multiplication by a constant $a$. That is for all $y\in Y$: $j(f(y))=a\cdot j(y)$ in $\mathbb{Z}/m\mathbb{Z}$ and hence the following diagram commutes:
@@
This result bridges the gap between arbitrary discrete maps and structured linear systems. Without loss of generality, we can identify elements of the finite set $Y$ of size $n$ with the set $\{0,1,\ldots,n-1\}$, which are elements of $\mathbb{Z}/n\mathbb{Z}$. The full formal proof of this result is in [5].
@@ 2 Main Results
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. The function $f$ is represented by an $n\times n$ adjacency matrix $A=A_{f}$, where the entry $a_{ij}=\delta_{f(i),j}$ and $\delta_{i,j}$ is the Kronecker delta. With this convention, each row of AA contains exactly one non-zero entry.
Lemma $2.$
2.1Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Then for all $i\in\{0,1,\ldots,n-1\}$ and $y\in\mathbb{Z}^{n}$
$(Ay)_{i}=y_{f(i)}.$
Proof.
The proof follows from the Definition 2.1.
$(Ay)_{i}=\sum_{j=0}^{n-1}a_{ij}y_{j}=\sum_{j=0}^{n-1}\delta_{f(i),j}y_{j}=y_{f(i)}$
@@
2.12.2Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function and $A=A_{f}$ be the adjacency matrix of the function $f$. Let $v\in\mathbb{Z}^{n}$, $y=\operatorname{adj}(xI-A)v$ and $m=\det(xI-A)$. Then for all $i\in\{0,1,\ldots,n-1\}$
$y_{f(i)}=xy_{i}-mv_{i}.$
@@
For adjugate matrix we have identity $(xI-A)\operatorname{adj}(xI-A)=\det(xI-A)I$. Therefore,
$mv=\det(xI-A)v=(xI-A)\operatorname{adj}(xI-A)v=(xI-A)y=xy-Ay.$
The final equality follows from the Lemma $2.$ ∎
@@
Let $M$ be an $n\times n$ matrix with polynomial entries $m_{ij}\in\mathbb{Z}[x]$. Then
$\deg(\det(M))\leq\sum_{i=0}^{n-1}\sum_{j=0}^{n-1}\deg(m_{ij}).$
@@
Lemma 2.5.
Let AA be an $n\times n$ matrix with integer entries. The characteristic matrix $\chi_{A}(x)=xI-A$ has determinant equal to the characteristic polynomial:
$\det(\chi_{A}(x))=\det(xI-A)$
For all $n\in N$, this polynomial is monic of degree $n$.
@@
This follows from the standard properties of the characteristic polynomial. ∎
Lemma 2.6.
2.4Let AA be an $n\times n$ matrix with integer entries and $\chi_{A}(x)=xI-A$ be its characteristic matrix. For $i\neq j$, the $(i,j)$ entry of $\operatorname{adj}(\chi_{A}(x))$ has degree at most n2n-2.
@@
The adjugate entry $\operatorname{adj}(\chi_{A}(x))_{ij}$ equals the determinant of the characteristic matrix with row $j$ and column $i$ removed from $\chi_{A}(x)$.
This submatrix has diagonal entries from $\chi_{A}(x)$ except at diagonal positions $i$ and $j$ (which are deleted). Since $\chi_{A}(x)$ has diagonal entries of degree 11 (from xIxI) and off-diagonal entries of degree 0 (from A-A), the submatrix has exactly n2n-2 diagonal entries of degree 11 and all other entries of degree 0.
By Lemma 2.4, the determinant has degree at most n2n-2. ∎
Lemma 2.7.
2.62.5Let $M=M(x)=\operatorname{adj}(xI-A)$ be the adjugate of the characteristic matrix xIAxI-A. Then the matrix entries $m_{ij}=p_{ij}(x)$ are polynomials in $x$ for all $i,j\in\{0,1,\ldots,n-1\}$ such that
  • $p_{ii}(x)$ is monic of degree n1n-1 for all $i\in\{0,1,\ldots,n-1\}$ and
  • $p_{ij}(x)$ has degree at most n2n-2 for all $i\neq j\in\{0,1,\ldots,n-1\}$.
@@
The diagonal entries of $\operatorname{adj}(xI-A)$ are characteristic polynomials of $(n-1)\times(n-1)$ submatrices, hence monic of degree n1n-1 by Lemma 2.5.
The off-diagonal case follows directly from Lemma 2.6. ∎
Definition 2.8.
For a polynomial $p(x)=\sum_{i=0}^{d}p_{i}x^{i}\in\mathbb{Z}[x]$, we define the coefficients bound:
$|p|=\sum_{i=0}^{d}|p_{i}|$
Lemma 2.9.
2.8If $p\in\mathbb{Z}[x]$ has positive leading coefficient, then for all integers $n\geq|p|$, we have $n>0$ and $p(n)>0$.
Proof.
Lemma is trivially true for $d=0$ so we can assume $d\geq 1$. Write $p(n)=an^{d}+r(n)$ where $a\geq 1$ is the leading coefficient of $p$, $d=\deg(p)$, and $\deg(r)<d$. For $n\geq|p|$:
$n\geq|p|\geq a\geq 1>0.$
Since $n\geq 1$, we have $|r(n)|\leq Bn^{d-1}$ where B=|p|aB=|p|-a. Therefore,
$p(n)=an^{d}+r(n)\geq an^{d}-Bn^{d-1}=n^{d-1}(an-B)\geq(a|p|-B)\geq|p|-B=a\geq 1>0.$
Lemma 2.10.
2.72.92.5Let $M=(m_{ij})=M(x)=\operatorname{adj}(xI-A)$ be the adjugate of the characteristic matrix xIAxI-A. Let $v=(1,2,\ldots,n)^{T}$ and $m=m(x)=\det(xI-A)$. Then for sufficiently large integer $x$:
$0<y_{0}<y_{1}<\cdots<y_{n-1}<m$
Proof.
The proof follows from Lemma 2.7 and Lemma 2.9. Let $y=Mv$. Then $y_{i}=\sum_{k=0}^{n-1}m_{ik}(k+1)$.
For each entry $y_{i}$, we express it as evaluation of a polynomial $p_{i}(x)=\sum_{k=0}^{n-1}m_{ik}(x)(k+1)\in\mathbb{Z}[x]$. By Lemma 2.7, the diagonal entry $m_{ii}$ is monic of degree n1n-1, while off-diagonal entries $m_{ik}$ (for $k\neq i$) have degree at most n2n-2. Therefore, the coefficient of $x^{n-1}$ in $p_{i}$ is $i+1>0$ (dominated by the $m_{ii}(i+1)$ term).
For the difference $p_{j}-p_{i}$ with $j>i$, the leading term comes from $(m_{jj}(j+1)-m_{ii}(i+1))$. Since both $m_{jj}$ and $m_{ii}$ are monic of degree n1n-1, the leading coefficient of $p_{j}-p_{i}$ is (j+1)(i+1)=ji>0(j+1)-(i+1)=j-i>0.
Similarly, for $p_{m}(x)=\det(xI-A)-p_{i}(x)$, since $\det(xI-A)$ is monic of degree $n$ (by Lemma 2.5) and $p_{i}$ has degree at most n1n-1, the leading coefficient is 1.
Since $p_{0}(x)$ has leading coefficient $0+1=1>0$, we have $y_{0}>0$ for sufficiently large $x$.
@@
Let $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ be any function. A linear representation of $f$ is an injective function $j:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/m\mathbb{Z}$ such that for all $i\in\{0,1,\ldots,n-1\}$,
@@
Theorem 2.12 (Main Theorem).
2.112.32.10Any finite function $f:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ has a linear representation.
@@
Let us define $m=\det(xI-A)$, $a=x$ and $j(i)=y_{i}$ for $i\in\{0,1,\ldots,n-1\}$, where $y=Mv$ with $v=(1,2,\ldots,n)^{T}$. By Lemma 2.10, we can choose $x$ large enough such that:
$0<y_{0}<y_{1}<\cdots<y_{n-1}<m$
Since $m$ is monic of degree $n$, we can choose $x$ such that $x<m(x)$. This shows that $j$ is injective and $a\in\{0,1,\ldots,m-1\}$.
Lemma 2.3 shows that $y_{f(i)}=xy_{i}-mv_{i}$ and hence $y_{f(i)}\equiv xy_{i}\pmod{m}$.
Therefore, $j$ is a linear representation of $f$ with modulus $m$ and constant $a$. ∎
@@
The adjacency matrix is:
@@
$xI-A_{f}=\begin{pmatrix}x-1&0&0\\ 0&x-1&0\\ 0&-1&x\end{pmatrix}$
@@
$\operatorname{adj}(xI-A_{f})=\begin{pmatrix}x(x-1)&0&0\\ 0&x(x-1)&0\\ 0&(x-1)&(x-1)^{2}\end{pmatrix}$
Using vector $v=(1,2,3)^{T}$, we get:
$y=\operatorname{adj}(xI-A_{f})\cdot v=\begin{pmatrix}x(x-1)\cdot 1\\ x(x-1)\cdot 2\\ (x-1)\cdot 2+(x-1)^{2}\cdot 3\end{pmatrix}=\begin{pmatrix}x^{2}-x\\ 2x^{2}-2x\\ (x-1)(3x-1)\end{pmatrix}=\begin{pmatrix}x^{2}-x\\ 2x^{2}-2x\\ 3x^{2}-4x+1\end{pmatrix}$
@@
$j(0)=12,\quad j(1)=24,\quad j(2)=33$
These values are strictly increasing and bounded by $m=36$, so $j$ is injective.
@@
The lemma states that $y_{f(i)}=xy_{i}-m\cdot v_{i}$, which we can rewrite as:
$j(f(i))\equiv xj(i)\pmod{m}$.
@@
$j(f(0))=j(0)=12$ $\equiv 4\cdot 12-36\cdot 1=48-36=12=4\cdot j(0)\pmod{36}\quad\checkmark$
$j(f(1))=j(1)=24$ $\equiv 4\cdot 24-36\cdot 2=96-72=24=4\cdot j(1)\pmod{36}\quad\checkmark$
$j(f(2))=j(1)=24$ $\equiv 4\cdot 33-36\cdot 3=132-108=24=4\cdot j(2)\pmod{36}\quad\checkmark$